Problem 6-5 Expected Return: Discrete Distribution A stocks return has the following distribution: Demand for the Companys
Answer #1

Expected return=Respective return*Respective probability

=(0.1*-45)+(0.2*-8)+(0.4*8)+(0.2*30)+(0.1*60)

=9.1%

probability Return probability*(Return-Expected Return)^2
0.1 -45 0.1*(-45-9.1)^2=292.681
0.2 -8 0.2*(-8-9.1)^2=58.482
0.4 8 0.4*(8-9.1)^2=0.484
0.2 30 0.2*(30-9.1)^2=87.362
0.1 60 0.1*(60-9.1)^2=259.081
Total=698.09%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(698.09)^(1/2)

=26.42%(Approx).